# Can SAEs reveal and mitigate racial biases of LLMs in healthcare?

### v1

@@ Abstract
LLMs are increasingly being used in healthcare. This promises to free physicians from drudgery, enabling better care to be delivered at scale. But the use of LLMs in this space also brings risks; for example, such models may worsen existing biases. How can we spot when LLMs are (spuriously) relying on patient race to inform predictions? In this work we assess the degree to which Sparse Autoencoders (SAEs) can reveal (and control) associations the model has made between race and stigmatizing concepts. We first identify SAE latents in gemma-2 models which appear to correlate with Black individuals. We find that this latent activates on reasonable input sequences (e.g., “African American”) but also problematic words like “incarceration”. We then show that we can use this latent to “steer” models to generate outputs about Black patients, and further that this can induce problematic associations in model outputs as a result. For example, activating the Black latent increases the risk assigned to the probability that a patient will become “belligerent”. We also find that even in this controlled setting in which we causally intervene to manipulate only patient race, elicited CoT reasoning strings do not communicate that race is a factor in the resulting assessments. We evaluate the degree to which such “steering” via latents might be useful for mitigating bias. We find that this offers improvements in simple settings, but is less successful for more realistic and complex clinical tasks. Overall, our results suggest that: SAEs may offer a useful tool in clinical applications of LLMs to identify problematic reliance on demographics, as compared to CoT explanations, which should not be trusted in such settings. But mitigating bias via SAE steering appears to be of marginal utility for realistic tasks  [^1] Code will be released at a later date..
@@ 1 Introduction
Specifically, using discharge summaries of patients who identify as Black or white, we train a linear probe on SAE activations to identify latents most predictive of race. We find that the latent with the highest estimated coefficient activates, intuitively, on mentions of Black identity. But it also fires on stigmatizing concepts like cocaine use and incarceration in clinical notes. To establish causality, we steer the model using this latent and observe that the model considers patients that are “more Black” to be at greater risk of getting belligerent. We then see if SAEs can be used to detect and mitigate racial bias in clinical generation tasks. For the simple task generating vignettes of patients with a clinical condition (Zack et al., 2024) we find that ablating the Black latent can reduce over-representation of Black patients when sampling cases for conditions such as cocaine abuse. However, when considering more complex tasks such as risk prediction based on clinical notes we observe that SAEs do not offer a reliable mechanism to mitigate racial bias.
@@
The key takeaways from this work are: SAE latent descriptions should be domain specific; Modern LLMs still have internalized problematic associations between race and input concepts in the high-stakes context of healthcare, and SAEs can reveal and characterize these in some cases, even where model reasoning (CoT) does not, and; SAEs can also be used to somewhat mitigate biases, but their utility on realistic clinical tasks relative to careful prompting remains an open question.
@@ 2 Locating race predictive latents
We follow Movva et al. (2025) and train a logistic regression probe with $\ell$1 regularization to predict race $\bm{y}$} from $\bm{Z}$}. Note that this task is not as trivial as looking for explicit mentions of race. Race is mentioned in only $4.3\%$ of our dataset.
We experiment with two models: gemma-2-2B-it and gemma-2-9B-it (Team et al., 2024), and use Gemmascope SAEs (Lieberum et al., 2024) of width $16\text{K}$ trained on the residual stream activations of the base model. Following prior work (et al., 2024; Bouzid et al., 2025) , we use the middle layer ($\ell=12$ for 2B and $\ell=20$ for 9B variant) SAEs for our analyses.
@@ 2.2 Results
Appendix Table 8 provides descriptions of the top-$5$ latents most predictive of race. The top latent in both models is about references to African-American ethnicity; we take these as the “Black latent” for the respective models. The AUROCs computed using the Black latent’s max-aggregated activations are $0.63$ and $0.72$ for gemma-2-2B and gemma-2-9B, suggesting that this single latent strongly correlates with patient race.
What tokens do these Black latents activate on? Is it simply literal occurrences of “Black” and “African-American”? If so, it would not be of much use from an interpretability perspective. To contextualize this in the clinical domain, we interpret the latents on discharge summaries and inspect the top-activating examples. Figures 1(a) shows these for gemma-2-2B. The latent indeed strongly activates on occurrences of “Black” and “African-American”. Further, it activates on conditions that are comparatively prevalent in the Black population, such as preterm labor (Manuck, 2017) and gestational hypertension (Ford, 2022). However, it also activates on tokens suggesting problematic implicit associations: incarceration, gunshot wounds, altercation with the police, and cocaine use (examples in red boxes). We see similar associations in gemma-2-9B (Figure 1(b)).
(a) gemma-2-2B
(b) gemma-2-9B
Figure 1: Max-activating examples of Black latents in clinical discharge summaries. The latents activate on mentions of Black identity, which is intuitive. But they also reveal problematic associations like activating on cocaine (examples boxed in red).
@@ 3 Steering with the Black latent
Does the Black latent merely reveal racial associations with input tokens, or does it (also) have a causal effect on the model’s output? To answer this we evaluate steering performance using the latent. Given that we observe the latent highly activate on discussions related to violence (altercation with the police, gunshot wounds, incarceration) in both models, we evaluate whether steering with the Black latent induces the model to view patients as violent.
@@
We use the brief hospital courses (BHC) dataset provided by Hegselmann et al. (2024) and sample $500$ BHCs for evaluation and $100$ for selecting $\alpha$. For the latter, we compute $\frac{\text{Positive rate}}{\text{Perplexity}}$ for a range of $\alpha$ values ($\alpha\in[0.01,5]$), where positive rate is the fraction of outputs with ‘Yes’ and perplexity is the output perplexity computed under Llama-3.1-8B (Dubey et al., 2024). We account for perplexity to ensure the interventions do not deteriorate text quality (see Appendix C for a plot of $\alpha$ vs. perplexity).
@@
2B $\uparrow$ $0.51$ 0.01-0.01 $1.0$ $0.80$
9B $\uparrow$ $0.80$ $0.09$ $0.78$ $1.0$
In Table 3, $\text{race}_{\text{Black}}$ reports the fraction of outputs that explicitly indicate the patient is Black (recall that we prompt the model to include this information). Steering is effective at assigning the patient race, indicating that these latents are not only input-centric but also mediate a causal effect on outputs. We also report $\Delta_{\text{Black}}$, which is the change in the positive rate after steering. We observe that increasing the “Black”-ness of a patient—that is, increasing the “Black” latent activation per Equation 1—increases the predicted risk of patient belligerence.
@@
Is model Chain-of-Thought (CoT) faithful to its internal reasoning when it relies on race? Above we showed that we can reliably “assign” race to a patient via steering and this causally induces meaningfully different predictions regarding their likelihood of becoming violent. Will CoT reveal this as a driving factor? To assess this we search for occurrences of race-related terms (such as ‘African’, ‘Black’, ‘racial’) in the model’s steered (CoT) outputs.
@@ 4.2 Clinical Tasks
Our goal here is not to completely remove the model’s ability to represent and/or factor race into its predictions. This would enforce demographic parity (Barocas et al., 2020), where the model’s positive rate is unaffected by race. Demographic parity can be problematic in the clinical domain as relevant clinical features associated with race may be ignored, introducing biases in another dimensions (Pfohl et al., 2019; Zhang et al., 2020). Instead, we are interested in detecting and mitigating reliance on race when irrelevant to the task, as well as providing transparency into when race is a driving factor for an output.
@@ 4.2.1 Tasks
Risk prediction
Prior work (Ahsan et al., 2024) has shown that LLMs can be effective in retrieving evidence for a suspected diagnosis from patient history. Given a patient note and a clinical condition, an LLM is prompted to determine if the patient is at risk of the condition based on the information present in the note. While conditions like gestational hypertension and uterine fibroids (Katon et al., 2023) are more prevalent in the Black population, this setting—where the model is tasked with drawing evidence for this specific patient from their history—rewards reasoning directly over the individual and detailed patient history; broad correlations informed by demographic categories should be less useful here. And if the model does use race to inform its output, we would like the model to state this explicitly in its explanation; but as shown in Section 3, the model can be unfaithful to its internal reasoning in this respect.
For patient notes, we use the brief hospital courses (BHC) dataset introduced in Section 3. To allow comparison with the vignette generation setup, we evaluate the same conditions considered in Section 4.1. For cocaine abuse, we sample BHCs with the term ‘cocaine’. To ensure a BHC does not contain discussions about cocaine abuse (which would make the task trivial), we filter BHCs with the term ‘cocaine abuse’.
@@
Prior studies (Lee et al., 2019; Morales & Yong, 2021) have shown that Black patients are less likely to be offered pain medication compared to white patients, even when they are experiencing comparable levels of pain. To investigate whether LLMs exhibit similar biases, Logé et al. (2021) introduced the Q-Pain dataset, comprising medical scenarios in which different gender and racial profiles can be substituted. In all cases, the patient should receive pain medication [^3] The dataset contains one case in which the patient does not need pain medication; this was provided for few-shot prompting. We ignore the particular case.. To study demographic bias, an LLM is prompted to determine if the patient should be given pain medication and one can measure whether the output changes when different demographic groups are substituted. Here we compare changes in outputs when race is surgically altered.
@@ 4.2.2 Approach
Our goal is to detect whether race affects an LLM’s output and, if so, if the effect can be mitigated using SAEs. We first identify causally important latents. We adopt the method of Marks et al. (2025) to identify causally important latents by approximating the effect ablating each latent has on the model output. Given an output metric $m$, the effect $E$ of a latent activation $z$ is
$E=\sum_{t}\big(m(x|\text{do}(z_{t}=0))-m(x)\big)$ (2)
where $x$ is the input and $z_{t}$ is the latent activation at token position $t$: This sums over the effects of intervening on latents at each token position. Here we are interested in differences between risk predictions made for Black patients as compared to other individuals. Concretely, we measure this as $m=\text{logit (``Yes'')}-\text{logit (``No'')}$ for a given $x$.
@@ 4.2.3 Results
Figure 3 shows $\Delta_{\text{logitdiff}}$ for all tasks. Prompting with an anti-bias statement significantly reduces $\Delta_{\text{logitdiff}}$ in four out of five tasks. For cocaine overdose, the model seems to over-correct and significantly shifts towards generating ‘Yes’ for white patients. Zero-ablating SAE race latents does not affect the output in three out of five tasks. It marginally reduces logit difference in risk prediction for uterine fibroids and gestational hypertension by $0.05$ and $0.03$ respectively. We also experiment with ablating race latents simultaneously in five layers (middle layer onwards) but see no improvement in performance (see Appendix D.3 for FLDD).
@@
Summary
Our results on the use of SAEs to recognize and mitigate implicit LLM racial bias in healthcare are mixed. On the somewhat contrived task of “vignette generation”, we offer promising findings. But on more realistic and complex tasks, the effect of ablating racial latents is minimal.
@@ 5 Related Work
6 Conclusions
Our results show that SAEs can reveal problematic associations about patients and race, and permit interventions that are effective in some settings. However, results on more complex clinical tasks are mixed. Perhaps representation of race in simple tasks is comparatively localized, so intervening on even a single latent can significantly affect model output; race maybe more dispersed and entangled with clinical concepts in more realistic and complex clinical tasks.
Limitations
@@ 6 Conclusions
Acknowledgements
This work was supported by a grant from Open Philanthropy.
@@ Appendix A Datasets
We use the dataset, ‘Medical Expert Annotations of Unsupported Facts in Doctor-Written and LLM-Generated Patient Summaries’, introduced by Hegselmann et al. (2024), licensed under The PhysioNet Credentialed Health Data License Version 1.5.0  [^5] https://physionet.org/content/ann-pt-summ/view-license/1.0.0/. The dataset is derived from MIMIC-IV-Note v2.2 database (Johnson et al., 2023) which includes $331,793$ deidentified free-text clinical notes from $145,915$ patients admitted to the Beth Israel Deaconess Medical Center in Boston, MA, USA. We use the MIMIC-IV-Note-Ext-DI-BHC subset, which contains Brief Hospital Courses (BHC)-summary pairs. We use the BHCs in the train-split (train.json).‘
We also use Q-Pain dataset (Logé et al., 2021) licensed under the Creative Commons Attribution-ShareAlike 4.0 International Public License footnotehttps://www.physionet.org/content/q-pain/view-license/1.0.0/

### v2

@@ Abstract
LLMs are increasingly being used in healthcare. This promises to free physicians from drudgery, enabling better care to be delivered at scale. But the use of LLMs in this space also brings risks; for example, such models may worsen existing biases. How can we spot when LLMs are (spuriously) relying on patient race to inform predictions? In this work we assess the degree to which Sparse Autoencoders (SAEs) can reveal (and control) associations the model has made between race and stigmatizing concepts. We first identify SAE latents in gemma-2 models that appear to correlate with Black individuals. We find that these latents activate on reasonable input sequences (e.g., “African American”) but also problematic words like “incarceration”. We then show that we can use this latent to “steer” models to generate outputs about Black patients, and further that this can induce problematic associations in model outputs as a result. For example, activating latents associated with Black individuals increases the risk assigned to the probability that a patient will become “belligerent”. We also find that even in this controlled setting where we causally intervene to manipulate only patient race, elicited CoT reasoning chains do not communicate race as a factor in the resulting assessments. We evaluate the degree to which such “steering” via latents might be useful for mitigating bias. We find that this offers improvements in simple settings, but is less successful for more realistic and complex clinical tasks. Overall, our results are mixed, and suggest that SAEs may offer a useful tool in clinical applications of LLMs to identify problematic reliance on demographics, as compared to CoT explanations (which should not be trusted in such settings). But mitigating bias via SAE steering may be only marginally effective for more realistic tasks.
@@ 1 Introduction
Specifically, using discharge summaries of patients who identify as Black or white, we train a linear probe on SAE activations to identify latents most predictive of race. We find that the latent with the highest estimated coefficient activates, intuitively, on mentions of Black identity. But it also fires on stigmatizing concepts like cocaine use and incarceration in clinical notes. To establish causality, we steer the model using this latent and observe that the model considers patients that are “more Black” to be at greater risk of becoming belligerent. We then see if SAEs can be used to detect and mitigate racial bias in clinical generation tasks. For the simple task of generating vignettes of patients with a clinical condition (Zack et al., 2024) we find that ablating the Black latent can reduce over-representation of Black patients when sampling cases for conditions such as cocaine abuse. However, when considering more complex tasks such as risk prediction based on clinical notes, we observe that SAEs do not offer a reliable mechanism to mitigate racial bias.
@@
The key takeaways from this work are: SAE latent descriptions should be domain specific; Modern LLMs still have internalized problematic associations between race and input concepts in the high-stakes context of healthcare, and SAEs can reveal and characterize these in some cases, even where model reasoning (CoT) does not, and; SAEs can also be used to somewhat mitigate biases, but their utility on realistic clinical tasks relative to careful prompting remains an open question. We release code at https://github.com/hibaahsan/sae_bias/.
@@ 2 Locating race predictive latents
We follow Movva et al. (2025) and train a logistic regression probe with $\ell$1 regularization to predict race $\bm{y}$} from $\bm{Z}$}. Note that this task is not as trivial as looking for explicit mentions of race: Race is mentioned in only $4.3\%$ of notes in our dataset.
We experiment with two models: gemma-2-2B-it and gemma-2-9B-it (Team et al., 2024), and use Gemmascope SAEs (Lieberum et al., 2024) of width $16\text{K}$ trained on the residual stream activations of the base model. Following prior work (et al., 2024; Bouzid et al., 2025) , we use the middle layer ($\ell=12$ for 2B and $\ell=20$ for 9B) SAEs for our analyses.
@@ 2.1 Reinterpreting latents using clinical text
(a) gemma-2-2B
(b) gemma-2-9B
Figure 1: Max-activating examples of Black latents in clinical discharge summaries. Highlight intensities reflect activation strength. The latents activate on mentions of Black identity and conditions that are relatively common in Black patients, which is intuitive. But they also reveal problematic associations (boxed in red) like activating on cocaine.
@@ 2.2 Results
Appendix Table 8 provides descriptions of the top-$5$ latents most predictive of race. The top latent in both models is about references to African-American ethnicity; we take these as the “Black latent” for the respective models. The AUROCs computed using the Black latent’s max-aggregated activations are $0.63$ and $0.72$ for gemma-2-2B and gemma-2-9B, suggesting that this single latent strongly correlates with Black identity (see Figures 4 and 5 for activation pattern on the general-domain corpus).
What tokens do these Black latents activate on? Is it simply literal occurrences of “Black” and “African-American”? If so, it would not be of much use from an interpretability perspective. To contextualize this in the clinical domain, we interpret the latents on discharge summaries and inspect the top-activating examples. Figure 1(a) shows these for gemma-2-2B. The latent indeed strongly activates on occurrences of “Black” and “African-American”. [^2] The stronger activation on the token “American” is because knowledge about a multi-token entity is often stored in its last token (Meng et al., 2022). Further, it activates on conditions that are comparatively prevalent in the Black population, such as preterm labor (Manuck, 2017) and gestational hypertension (Ford, 2022). However, it also activates on tokens that suggest problematic implicit associations: Incarceration, gunshot wounds, altercation with the police, and cocaine use (examples in red boxes). We see similar associations in gemma-2-9B (Figure 1(b)). This finding that such latents that can reveal problematic associations generalizes to other model families and bigger models—see Appendix B.1 for examples for gpt-oss-20b (OpenAI et al., 2025).
@@ 3 Steering with the Black latent
Does the Black latent merely reveal racial associations with input tokens, or does it (also) have a causal effect on the model’s output? To answer this we evaluate steering performance using the latent. Given that we observe the latent highly activate on discussions related to violence (altercation with the police, gunshot wounds, incarceration) in both models, we evaluate whether steering with the Black latent induces the model to view patients as (potentially) violent.
@@
[…]
2B $\uparrow$ $0.51$ 0.01-0.01 $1.0$ $0.80$
9B $\uparrow$ $0.80$ $0.09$ $0.78$ $1.0$
@@
In Table 2, $\text{race}_{\text{Black}}$ reports the fraction of outputs that explicitly indicate the patient is Black (recall that we prompt the model to include this information). Steering is effective at assigning the patient race, indicating that these latents are not only input-centric but also mediate a causal effect on outputs. We also report $\Delta_{\text{Black}}$, which is the change in the positive rate after steering. We observe that increasing the “Black”-ness of a patient—that is, increasing the “Black” latent activation per Equation 1—increases the predicted risk of patient belligerence. Table 3 shows a sample output from gemma-2-2b-it before and after steering.
@@
Is model Chain-of-Thought (CoT) faithful to its internal reasoning when it relies on race? Above we showed that we can reliably “assign” race to a patient via steering and this causally induces meaningfully different predictions regarding their likelihood of becoming violent. Will CoT mention this as a factor? To assess this we search for occurrences of race-related terms (such as ‘African’, ‘Black’, ‘racial’) in the model’s steered (CoT) outputs.
@@ 4.2 Clinical Tasks
Our goal here is not to completely remove the model’s ability to represent and/or factor race into its predictions. This would enforce demographic parity (Barocas et al., 2020), where the model’s positive rate is unaffected by race. Demographic parity can be problematic in the clinical domain as relevant clinical features associated with race may be ignored, introducing biases in another dimensions (Pfohl et al., 2019; Zhang et al., 2020). Instead, we are interested in detecting and mitigating reliance on race when irrelevant to the task.
@@ 4.2.1 Tasks
Diagnosis evidence
Prior work (Ahsan et al., 2024) has shown that LLMs can be effective in retrieving evidence for a suspected diagnosis from patient history. Given a patient note and a clinical condition, an LLM is prompted to determine if the patient is at risk of the condition based on the information present in the note. While conditions like gestational hypertension and uterine fibroids (Katon et al., 2023) are more prevalent in the Black population, this setting—where the model is tasked with drawing evidence for this specific patient only from their history—rewards reasoning directly over the individual and detailed patient history; broad correlations informed by demographic population prevalence is not desirable here. And if the model does use race to inform its output, we would like the model to state this explicitly in its explanation; but as shown in Section 3, the model can be unfaithful to its internal reasoning in this respect.
For patient notes, we use the brief hospital courses (BHC) dataset introduced in Section 3. To allow comparison with the vignette generation setup, we evaluate the same conditions considered in Section 4.1. For cocaine abuse, we sample BHCs of male patients with the term ‘cocaine’. To ensure a BHC does not contain discussions about cocaine abuse (which would make the task trivial), we filter BHCs with the term ‘cocaine abuse’.
@@
Prior studies (Lee et al., 2019; Morales & Yong, 2021) have shown that Black patients are less likely to be offered pain medication compared to white patients, even when they are experiencing comparable levels of pain. To investigate whether LLMs exhibit similar biases, Logé et al. (2021) introduced the Q-Pain dataset, comprising medical scenarios in which different gender and racial profiles can be substituted. In all cases, the patient should receive pain medication. [^4] The dataset contains one case in which the patient does not need pain medication; this was provided for few-shot prompting. We ignore the particular case. To study demographic bias, an LLM is prompted to determine if the patient should be given pain medication and one can measure whether the output changes when different demographic groups are substituted. Here we compare changes in outputs when race is surgically altered.
@@ 4.2.2 Approach
Our goal is to detect whether race affects an LLM’s output and, if so, if the effect can be mitigated using SAEs. We first compute the causal effect of latents. We adopt the method of Marks et al. (2025) and approximate the effect ablating each latent has on the model output. Given an output metric $m$, the effect $E$ of a latent activation $z$ is
$E=\sum_{t}\big(m(x)-m(x|\text{do}(z_{t}=0))\big)$ (2)
where $x$ is the input and $z_{t}$ is the latent activation at token position $t$: This sums over the effects of intervening on latents at each token position. Here we are interested in differences between predictions made for Black patients as compared to other individuals. Concretely, we measure this as $m=\text{logit (``Yes'')}-\text{logit (``No'')}$ for a given $x$.
@@ 4.2.3 Results
Figure 3 shows $\Delta_{\text{logitdiff}}$ for all tasks. Prompting with an anti-bias statement significantly reduces $\Delta_{\text{logitdiff}}$ in four out of five tasks. For cocaine overdose, the model seems to over-correct and significantly shifts towards generating ‘Yes’ for white patients. Zero-ablating SAE race latents does not affect the output in three out of five tasks. It marginally reduces logit difference in risk prediction for uterine fibroids and gestational hypertension by $0.05$ and $0.03$ respectively.
We also experiment with ablating race latents simultaneously in five layers (middle layer onwards) but see no improvement in performance (see Appendix D.3 for FLDD).
@@
5 Discussion
Our results show that SAEs can reveal problematic associations about patients and race, and permit interventions that are effective in some settings. On the somewhat contrived task of “vignette generation”, we report promising findings. But on more realistic and complex tasks, the effect of ablating racial latents is minimal. Perhaps representation of race in simple tasks is comparatively localized, so intervening on even a single latent can significantly affect model output; race maybe more dispersed and entangled with clinical concepts in more realistic and complex clinical tasks.
@@ 6 Related Work
7 Limitations
@@
8 Ethics
In Section 3, we show how model internals can be manipulated to induce harmful behavior. This exercise was performed to highlight problematic racial associations in LLMs. We caution against using such interventions intentionally to cause harm.
9 Reproducibility
We release code at https://github.com/hibaahsan/sae_bias/. We conduct experiments with HuggingFace implementations of models and use NNsight (Fiotto-Kaufman et al., 2024) to perform interventions. We use two NVIDIA H200 GPUs. We describe the datasets we used in Sections 3, 4.2, and in Appendix A.
10 Acknowledgements
This work was supported by Coefficient Giving. We thank David Bau, Can Rager, and Koyena Pal for their thoughtful feedback. We thank Caden Juang for providing the activation visualization interface.
This work used DeltaAI at NCSA through allocation CIS251008 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S. National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
@@ Appendix A Datasets
We use the dataset, “Medical Expert Annotations of Unsupported Facts in Doctor-Written and LLM-Generated Patient Summaries”, introduced by Hegselmann et al. (2024), licensed under The PhysioNet Credentialed Health Data License Version 1.5.0  [^6] https://physionet.org/content/ann-pt-summ/view-license/1.0.0/. The dataset is derived from MIMIC-IV-Note v2.2 database (Johnson et al., 2023) which includes $331,793$ deidentified free-text clinical notes from $145,915$ patients admitted to the Beth Israel Deaconess Medical Center in Boston, MA, USA. We use the MIMIC-IV-Note-Ext-DI-BHC subset, which contains Brief Hospital Courses (BHC)-summary pairs. We use the BHCs in the train-split (train.json).‘
We also use Q-Pain dataset (Logé et al., 2021) licensed under the Creative Commons Attribution-ShareAlike 4.0 International Public License  [^7] https://www.physionet.org/content/q-pain/view-license/1.0.0/
@@ Appendix B Race predictive latents
B.1 Black latent in gpt-oss-20b
Here we report findings using gpt-oss-20b. We used the open-source SAE corresponding to the middle layer (as we did with Gemma models) available on Neuronpedia and HuggingFace  [^8] https://huggingface.co/andyrdt/saes-gpt-oss-20b/tree/main/resid_post_layer_11/trainer_0. We found a latent that activates on mentions of Black population but also on stigmatizing concepts similar to those in Gemma. Figure 6 shows max-activating examples, similar to Figures 1.
Figure 6: Max-activating examples of Black latent in clinical discharge summaries using gpt-oss-20B. We see a similar pattern as observed in gemma-2 models—the Black latent activates on mentions of Black identity but also on stigmatizing associations.